0.5m + 6.4 = 4.9 -0.1m
Answer by coyote(32)

0.5m + 6.4 = 4.9 -0.1m
Add .1m to both sides
.6m + 6.4 = 4.9
Subtract 6.4 from both sides
.6m = -1.5
Divide both sides by .6
m = -2.5
-----
To check, substitute all "m"s with -2.5
.5(-2.5) + 6.4 = 4.9 - .1(-2.5)
-1.25 + 6.4 = 4.9 - (-.25)
5.15 = 5.15